THE DIRECT OBSERVATION AND ANALYSIS OF PSEUDOROTATION IN 1,3-DIOXOLANE

\underline{DMITRY G. MELNIK}, TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210; FRANK C. DE LUCIA, The Ohio State University, Dept. of Physics, Microwave Laboratory, 174 W. 18th Avenue, Columbus OH 43210.

A number of transitions identified as belonging to the rotational structure of n = 0 → n = 1, n = 2 → n = 3, and n = 0 → n = 3 pseudorotational (PR) bands in 1,3-dioxolane were observed in the direct absorption experiment in the range of 153-364 GHz using the pulsed jet FASSST spectrometer.^{a} The newly observed transitions have been analyzed together with the data available from microwave measurements.^{b} The data were globally fit to a Hamiltonian similar to that which has been used in the analysis of tetrahydrofuran,^{c} and the molecular constants of 1,3-dioxolane in the n = 0...3 pseudorotational states were determined. The frequencies of the newly observed PR bands, n = 2 → n = 3 and n = 0 → n = 3, along with the results of the microwave^{b} and the low resolution FIR studies^{d} were used in a global analysis to derive the potential barrier to pseudorotation in the molecule. Ab initio calculations were also used to describe the potential surface along the pseudorotational path. The results of the modeling and comparisons between ab initio and empirical surfaces will be discussed.

aD.Melnik, S.Gopalakrishnan, T.A.Miller, S.P.Belov and F.C.De Lucia, J.Chem.Phys., 114, 6100 (2001)

bP.A.Baron and D.O.Harris, J.Mol.Spec., 49, 70 (1974)

cD.Melnik, S.Gopalakrishnan, T.A.Miller and F.C.De Lucia, J.Chem.Phys., 118, 3589 (2003)

dJ.A.Greenhouse and H.L.Strauss, J.Chem.Phys., 50, 124 (1969)